OpenStudy (anonymous):

Please help! Fan+Medal!!! Questions in comment! :D If you help me, Ill also give a nice testimony!

OpenStudy (anonymous):

1. Translate the sentence into an equation. Six less than a number is 14. A. 6 – n = 14 B. 6n = 14 C. n – 6 = 14 D. n = 14 – 6

OpenStudy (anonymous):

n - 6 = 14, c.

OpenStudy (anonymous):

this is because it says 6 less than n.

OpenStudy (anonymous):

Theres another one

OpenStudy (anonymous):

four more than 6 times a number is 34 6n + 4 = 34

OpenStudy (anonymous):

the difference of 4 times a number and 6 is 34 4n - 6 = 34

OpenStudy (anonymous):

six more than 4 times a number is 34 6 + 4n = 34 it could also be 4n + 6 = 34, i was not looking at how the answer looked.

OpenStudy (anonymous):

the sum of four times an number and 6 is 34 6 + 4n = 34 it could also be 4n + 6 = 34, i was not looking at how the answer looked. yes, they had the same answer. weird.

OpenStudy (anonymous):

done, @Korey_Minter !

OpenStudy (anonymous):

:D ALRIGHT 3 MORE

OpenStudy (anonymous):

i can try. these are fairly easy, though. so i can.

OpenStudy (anonymous):

wait im confused to is the second one i asked like what is 1.2.3.4.? wich letter is it

OpenStudy (anonymous):

1 is C 2 is B 3 is A 4 is A

OpenStudy (anonymous):

that helps?

OpenStudy (anonymous):

Sue has four $20 bills in her wallet. She needs $200 in cash to buy a new MP3 player. How many more $20 bills does Sue need to buy the MP3 player? Sue needs more $20 bills.

OpenStudy (anonymous):

Thats the last one! :D @silverr

OpenStudy (anonymous):

one second.

OpenStudy (anonymous):

nevermind i got it! :D its 6 :D

OpenStudy (anonymous):

that's not correct, @Korey_Minter ^^ 200 - 20 = ? ? / 20 = how many more she needs. can you do that for me?

OpenStudy (anonymous):

i put in 6 and it was right though

OpenStudy (anonymous):

really? hmm i was off. lol.